# Human Insights Driven Latent Space for Different Driving Perspectives: A Unified Encoder for Efficient Multi-Task Inference

### v1

@@ V Discussion and Conclusion
In terms of conclusion, we presented in this work a unified encoder achieved through multi-task learning for autonomous driving, leveraging human-like visual perception needed for navigation. By integrating knowledge from critical computer vision tasks such as depth, pose, 3D scene flow estimation, and various segmentation tasks into a single encoder, our approach allows efficient and compact multi-task inference. We addressed key challenges in multi-task training in ensuring the efficiency on each task. First, we introduce a novel multi-scale pose decoder to better estimate the relative pose between frames and improve the depth performance on scenes with few dynamic objects such as KITTI. Second, we propose to apply knowledge distillation from a multi-encoder model pretrained on the same tasks to stabilize the training process. Our experimental results demonstrate that the performance on individual tasks is competitive with state-of-the-art methods within a single encoder. Finally, when directly leveraging latent space from the unified encoder, such architecture is able to achieve coherent steering angle estimation by incorporating diverse encoded visual information for robust autonomous navigation.

### v2

@@ VI Conclusion
We addressed key challenges associated with multi-task training to ensure the effectiveness of each individual task. First, we introduced a novel multi-scale pose decoder, which enhances relative pose estimation between frames and improves depth performance, especially in dynamic scenes such as those in the KITTI dataset. Second, we employed knowledge distillation from a multi-encoder teacher model pretrained on the same tasks to stabilize the training process. Our experimental results demonstrate that the unified encoder achieves performance on par with state-of-the-art methods for individual tasks. Moreover, when directly leveraging the latent space from the unified encoder, the architecture is capable of coherent steering angle estimation, benefiting from the diverse encoded visual information. This highlights the potential of multi-task learning to advance robust and interpretable autonomous navigation systems.